# NICE: To Optimize In-Context Examples or Not?

### v1

@@ Abstract
Recent works have shown that large language models (LLMs) work remarkably well on a wide range of tasks through in-context learning and optimization of in-context examples (ICE). However, most of these studies assume either a fixed or no instruction provided in the prompt, leading to the apparent consensus that the optimization of in-context examples is critical for better performance. We challenge this consensus for instruction-tuned LLMs by investigating the necessity of optimizing in-context examples when task-specific instructions are provided, and find that there are tasks for which various ways of optimizing in-context examples yield diminishing returns. We introduce a task-specific metric called Normalized Invariability to Choice of Examples (NICE) that quantifies the learnability of tasks from a given instruction, and provides a heuristic that helps decide whether to optimize for instructions or ICE for any new task. On a wide range of tasks and a systematically created instruction set with gradually added details, we validate our hypothesis empirically by computing NICE with query-dependent bins of examples, comparing different instructions with ICE selection methods, and performing label perturbation experiments. We conclude that tasks can be divided into two broad classes based on the NICE metric, where the returns on ICE optimization follow predictable trends when instructions are provided in the prompt.
@@ 1 Introduction
The scaling up of large language models (LLMs) in terms of model size, training data, and compute in recent times has led researchers to notice their emergent in-context learning ability across a wide range of tasks (Brown et al., 2020; Radford et al., 2019; Kaplan et al., 2020). The in-context learning paradigm studies the ability to learn a new task based on an input context augmented with a few example demonstrations. The selection of good in-context examples (ICE) from a large pool of candidate training examples is a non-trivial optimization problem over a large search space (Liu et al., 2021).
@@
The emergence of in-context learning ability in LLMs has motivated a growing body of literature on how to select the ICE to be added to an LLM’s input prompt (Liu et al., 2021; Lu et al., 2021; Zhang et al., 2022b; Qin et al., 2023). These papers propose various ICE selection techniques that can significantly improve LLMs’ performance across a range of tasks, leading to the apparent consensus in literature that optimizing ICE is critical for better in-context learning performance. Thus, it appears that improving the optimization of ICE is the key to better in-context learning using LLMs.
@@
However, with the advent of instruction-tuning Wei et al. (2021); Chung et al. (2022); Ouyang et al. (2022), LLMs are also capable of following natural language instructions, making it possible to solve a task in a zero-shot manner, i.e., learning with just the instruction and no ICE. A natural question that follows is how instructions and ICE interact and whether they can be jointly optimized. Would optimization of ICE still matter after being given a descriptive or detailed task instruction? And more practically, given a task, what is better—optimizing the instruction or optimizing the ICE?
@@
Using state-of-the-art LLMs, our main finding is that the returns on ICE optimization diminish as more detailed instructions are added to the input prompt. We find that a detailed instruction with randomly chosen ICE surpasses or matches the accuracy of a prompt with ICE selected using criteria from past work Liu et al. (2021); Levy et al. (2022); Rubin et al. (2021); Sorensen et al. (2022), across diverse classification tasks such as SST-2 and SST-5 for 2 and 5-way sentiment classification respectively Socher et al. (2013), MNLI for natural language inference Williams et al. (2017), TREC for 6-way question classification Li and Roth (2002), MedQA Jin et al. (2021) and FinQA for domain-specific question-answering Chen et al. (2021). Moreover, given a detailed instruction, even using ICE with incorrect labels provides the same accuracy as using the correct ICE on most tasks. On the other hand, there exists a different class of tasks, especially generative tasks such as MTOP Li et al. (2020) for semantic parsing and Break Wolfson et al. (2020) for Question Decomposed Meaning Representation (QDMR) generation, where ICE selection actually matters: a prompt containing the best chosen ICE outperforms a prompt containing a descriptive instruction with random ICE.
@@
Therefore, contrary to prior work, ICE optimization per query may not always be the most effective strategy for increasing task performance, especially as task instructions are much simpler to write (e.g., using our proposed template). To characterize the tasks where ICE optimization may not matter, we introduce a metric that measures the invariability of an LLM’s task performance to ICE from different bins based on a query-dependent scoring function (e.g., cosine similarity). The metric, called Normalized Invariability to Choice of Examples (NICE), partitions candidate training examples into bins based on the above scoring function, and compares the performance of the best bin for picking random ICE with the mean performance across all bins. We show that this simple metric can effectively capture the dependence of tasks on ICE optimization. Tasks such as MTOP and BREAK score low on NICE metric (<0.4) whereas all the classification tasks score high (>0.85). Further, consistent with our results above, the NICE score increases as more detailed instructions are introduced in the prompt.
Figure 1: An overview of prompt optimization by improving the instruction and ICE following the NICE metric based for any new task following our hypothesis.
@@
As a practical implication, our work provides a method to decide the best option for improving an LLM’s task accuracy: ICE optimization or improving the instruction (see Figure 1). Given a task, we provide a template for creating a detailed instruction. Then, given the task and an instruction from the template, we provide an efficient procedure to compute the NICE score. If the NICE score is high, then it may be more effective to try improving the instruction; ICE optimization may not add much value. However, for tasks where it is low, ICE optimization will be effective.
To summarize, we make the following key contributions:
@@
  • Contrary to prior work, we show that ICE optimization does not always lead to performance improvement for LLMs. Across a range of tasks, we find that the performance of prompts improves with better-quality instructions, and this improvement remains constant even when true ICE are replaced with ICE with incorrect labels.
@@
  • For practical usage, we introduce a task-specific metric called NICE that can predict which tasks will benefit from ICE optimization.
@@
  • For tasks with high NICE score, we find that detailed instructions with random ICE obtain comparable or better performance than prompts with optimized ICE. Conversely, for tasks with low NICE score, prompts with optimized ICE obtain better performance.
In this study, we consider a prompt to contain both the instruction and the in-context examples, and compare the effectiveness of these components on a number of tasks. Previous works Min et al. (2022a); Wei et al. (2023) have argued that LLMs benefit from ICE even though they may not actually learn input-label mappings from ground truth examples. We find that the benefit of ICE changes with detailed instruction. There is a large class of tasks for which ICE optimization gives diminishing returns. Specifically, with the addition of a reasonable instruction in the prompt, dependence on ICE reduces significantly.
@@
2 Related Works
Several related works have studied the optimization of both instructions and ICE for LLMs on various classification and generation tasks.
2.1 In-context Example Selection Methods
Learning-free selection methods: Liu et al. (2021) propose a kNN-based method for selecting the top-$k$ candidates ranked in the reverse order of their cosine-similarity with the query based on the assumption that in-context examples with high similarity in the embedding space would help the most on downstream tasks. Sorensen et al. (2022) use a mutual-information-based metric for comparing two prompts and select the one that maximizes mutual information with the query. Gonen et al. (2022) attempt to select in-context examples with low perplexity, while Levy et al. (2022) show that diverse demonstrations improve the compositional generalization ability of LLMs.
Learning-based selection methods: Rubin et al. (2021) propose a two-stage retrieval method called Efficient Prompt Retrieval (EPR) to select ICE. For a specific input instance, they utilize a BM-25-based retriever to select candidate examples and subsequently employ a retriever trained through contrastive learning. Li et al. (2023) further enhance EPR by proposing a Unified Demonstration Retriever (UDR) across different tasks. Lu et al. (2022) propose a policy-gradient method for selecting in-context examples using a specific policy formulation for a given query. Ye et al. (2023) retrieve the complete set of demonstrations rather than extracting individual ones in order to capture inter-demonstration relationships. They used contrastive learning to train a Determinantal Point Process (DPP)-based retriever to match the scores of the language model output, obtaining the most likely demonstration set through maximum a-posteriori during inference.
2.2 Jointly Optimizing In-context Examples and Instructions
Some works (Zhou et al., 2022; Pryzant et al., 2023) use an LLM to improve the instruction alone but do not consider examples. Zhang et al. (2022a) introduce TEMPERA, an RL-based optimization scheme using discrete instruction and example sets. However, the search space for unconstrained joint optimization becomes exponentially large for large instruction and example spaces. While these works showcase the importance of optimizing both instructions and examples for various tasks, they do not perform a comprehensive comparison of the two. Our objective instead is to compare the two methods of task-learning for various tasks and evaluate how learning from instructions versus examples varies across tasks. We empirically show that the choice of optimal task-learning paradigm depends on the task, and we suggest a method to analyze any given task to gain some traction on which of these optimizations is expected to benefit model performance.
3 Formulating the Task-specific Metric
We propose a task-specific metric designed to categorize language tasks into two types based on their learnability from instructions and in-context examples. First, we employ a grouping function $f$ to partition the training set of candidate examples into multiple bins based on the query. Subsequently, for each bin, we compute the model’s performance by selecting random ICE from that bin for a set of sampled queries to get a score for each bin denoted by:
@@ 3 Formulating the Task-specific Metric
where $Y_{LM}$ is the performance measure (higher is better) of the language model, $\mathcal{D}$ is the set of test queries, I is the given instruction and $b^{x}_{j}\in\mathcal{B}$ is the $j$-th bin for query $x$. Next, we look at a few intuitive requirements from such a metric and finally define the NICE metric in Subsection 3.2.
@@ 3.1 Metric Requirements
Here we outline some intuitive requirements and considerations while formulating a metric that captures the non-benefit of ICE optimization for a given task and a given instruction:
  1. 1.
@@
  3. 3.
    Property 3: If $S(j,\text{I})\approx\max\limits_{j}S(j,\text{I})$ for all $j$, i.e., the performance of the model is invariant to the bin from which the in-context examples are chosen, the metric should be close to $1$.
  4. 4.
    Property 4: Let $b_{\text{max}}=\arg\max_{j}S(j,\text{I})$ be the best bin. If, for all $j\neq b_{\text{max}}$, $S(j,\text{I})\ll S(b_{\text{max}},\text{I})$, this represents a strictly retrieval task where the budget needs to be spent on in-context example optimization. In this scenario, the metric should approach 00.
Since API calls to LLMs are both costly and time consuming, another requirement is that the metric should be computable much more efficiently than typical ways to optimize over the search space of all candidate examples.
To compare different instructions, we can define a partial order on the set of instructions as $I\preceq I^{\prime}$ (in words, $I$ is Pareto-dominated by another instruction $I^{\prime}$) if $S(j,I)\leq S(j,I^{\prime})$, $\forall j$. One can prove (under simple assumptions) that if there exists an $I^{*}$ that is not Pareto-dominated by any other $I^{\prime}$, then it must be a maximizer of $\sum_{j}w_{j}S(j,I)$ for some non-negative weights $w_{j}$’s (Harsanyi, 1955). The weights $w_{j}$’s are typically uniform or chosen by domain experts (Harsanyi, 1955; Gale, 1960). For simplicity, we use uniform weights to define our metric. In the next section, we introduce the NICE metric, and in Appendix A.2, we prove that it follows all these requirements, and discuss additional benefits of its non-uniform variants.

### v3

@@ Abstract
Recent work shows that in-context learning and optimization of in-context examples (ICE) can significantly improve the accuracy of large language models (LLMs) on a wide range of tasks, leading to an apparent consensus that ICE optimization is crucial for better performance. However, most of these studies assume a fixed or no instruction provided in the prompt. We challenge this consensus by investigating the necessity of optimizing ICE when task-specific instructions are provided and find that there are many tasks for which it yields diminishing returns. In particular, using a diverse set of tasks and a systematically created instruction set with gradually added details, we find that as the prompt instruction becomes more detailed, the returns on ICE optimization diminish. To characterize this behavior, we introduce a task-specific metric called Normalized Invariability to Choice of Examples (NICE) that quantifies the learnability of tasks from a given instruction, and provides a heuristic to help decide whether to optimize instructions or ICE for a new task. Given a task, the proposed metric can reliably predict the utility of optimizing ICE compared to using random ICE. Our code is available at https://github.com/microsoft/nice-icl.
[…]
[backgroundcolor=gray!10, linecolor=black, outerlinewidth=0.5pt, roundcorner=1mm, skipabove=skipbelow=font=, ]promptbox
NICE: To Optimize In-Context Examples or Not?
Pragya Srivastavathanks:   These authors contributed equally.,   Satvik Golechha11footnotemark: 1,   Amit Deshpande,   Amit Sharma Microsoft Research, India {srivastava.pragya0805,zsatvik}@gmail.com, {amitdesh,amshar}@microsoft.com
@@ 1 Introduction
With the scaling up of large language models (LLMs) in terms of model size and training data, recent work demonstrates their emergent in-context learning ability across a wide range of tasks (Brown et al., 2020; Radford et al., 2019; Kaplan et al., 2020). Specifically, the in-context learning (ICL) paradigm studies ability of an LLM to perform a new task given a few example demonstrations. The selection of good in-context examples (ICE) from a large pool of candidates is a non-trivial optimization problem over a large search space (Liu et al., 2022).
@@
In-context learning ability in LLMs has motivated a growing body of literature on how to select the ICE to be added to an LLM’s input prompt (Liu et al., 2022; Lu et al., 2022; Zhang et al., 2022b; Qin et al., 2023). These works propose various ICE selection techniques and demonstrate significant improvements in LLM performance across a range of tasks, leading to an apparent consensus in the literature that ICE optimization is critical for better ICL performance of LLMs.
@@
However, with the advent of instruction-tuning Wei et al. (2021); Chung et al. (2024); Ouyang et al. (2022), LLMs are expected to follow natural language instructions, making it possible to solve a task in a zero-shot manner, i.e., learning with just the instruction and no ICE. As a result, a typical prompt to an LLM contains both the instruction and ICE. A natural question that follows is how instructions and ICE interact and whether they can be jointly optimized. Would ICE optimization still matter once a detailed task instruction is given? More practically, given a task, is optimizing the instruction more effective or optimizing ICE?
@@
Using state-of-the-art LLMs, our main finding is that returns on ICE optimization diminish as more detailed instructions are added to the input prompt. We find that a detailed instruction with randomly chosen ICE surpasses or matches the accuracy of a prompt with ICE selected using criteria from past work Liu et al. (2022); Levy et al. (2023); Rubin et al. (2021); Sorensen et al. (2022) across diverse tasks. These tasks include SST-2 and SST-5 for 2 and 5-way sentiment classification respectively Socher et al. (2013), MNLI for natural language inference Williams et al. (2017), TREC for 6-way question classification Li and Roth (2002), MedQA Jin et al. (2021) and FinQA for domain-specific question-answering Chen et al. (2021). Moreover, given a detailed instruction, even using ICE with incorrect labels yields the same accuracy as using correct labels on most tasks. On the other hand, there exists a different class of tasks, especially generative tasks that expect the output to follow specific schema, where ICE selection matters. For example, for tasks such as NL2BASH for bash command generation  Lin et al. (2018), MTOP Li et al. (2021) for semantic parsing, and Break Wolfson et al. (2020) for Question Decomposed Meaning Representation (QDMR) generation, a prompt with chosen examples outperforms a prompt containing a descriptive instruction with random ICE.
@@
Therefore, contrary to prior work, we argue that ICE optimization per query may not always be the most effective strategy for improving performance, especially if the task instructions are much simpler to write (e.g., using our proposed template in Section 3.1). To characterize the tasks where ICE optimization may or may not matter, we introduce a metric called Normalized Invariability to Choice of Examples (NICE) that measures, for a given task and an instruction, the invariability of an LLM’s task performance to ICE. We partition the candidate pool of examples into a relatively small number of query-dependent bins (e.g., based on cosine similarity of examples to the given query), and compute a bin-wise score given by the task performance of an LLM when the given task instruction is augmented with random ICE picked only from a particular bin. Our NICE metric compares the maximum score against the average score over all bins. We show that this simple metric can effectively capture the dependence of tasks on ICE optimization. Schema-based generative tasks such as MTOP and BREAK have low NICE values (<0.4) whereas all other tasks have high NICE values (>0.85). Further, the NICE value increases when we include more detailed task instructions, and suggests diminishing returns of ICE optimization for better instructions.
Figure 1: Prompt optimization using the NICE score. Given a task and a template instruction, the NICE score provides a heuristic to decide whether to improve the instruction (high NICE) or the in-context examples (low NICE) to optimize task performance.
@@
As a practical implication, our work provides a method to decide the best option for improving an LLM’s task accuracy: ICE optimization or improving the instruction (see Fig. 1). To this end, we propose a template for creating a detailed instruction for any task and provide an efficient procedure to compute the NICE metric. Given the task and a template instruction, if the NICE value is high, then it is more effective to improve the instruction and random ICE suffice. Otherwise, if the NICE value is low, it is more effective to do ICE optimization.
To summarize, our main contributions are:
@@
  • Contrary to prior work, we show that ICE optimization does not always lead to performance improvement for LLMs. Across a range of tasks, we find that the performance of prompts improves with more detailed instructions, and this improvement remains constant even when the true labels in ICE are replaced with random labels.
@@
  • We introduce a practical, task-specific metric called NICE that helps predict if a given task and an instruction will benefit from ICE optimization.
@@
  • For tasks with high NICE value, we find that detailed instructions with random ICE obtain comparable or better performance than prompts with optimized ICE. On the other hand, for tasks with low NICE value, optimized ICE lead to better performance.
2 Related Work
2.1 ICE Selection Methods
Prior work on ICE selection can be divided into learning-free and learning-based methods.
Learning-free methods: Liu et al. (2022) propose a kNN-based method for selecting the top-$k$ candidates based on their cosine-similarity with the query with the assumption that ICE with high similarity in the embedding space would help the most on downstream tasks. Sorensen et al. (2022) compare two in-context examples and select the one that maximizes the mutual information with the query. Gonen et al. (2023) attempt to select ICE with low perplexity, while Levy et al. (2023) show that diversity in demonstrations improves the compositional generalization ability of LLMs.
Learning-based methods: Rubin et al. (2022) propose a two-stage retrieval method to select ICE. For a given input, they utilize a BM-25-based retriever to select candidate examples and then employ a retriever trained through contrastive learning to pick ICE. Li et al. (2023) enhance it by proposing a Unified Demonstration Retriever (UDR) across different tasks. Lu et al. (2023) propose a policy-gradient method for selecting ICE using a specific policy formulation for a given query. Ye et al. (2023) retrieve a set of demonstrations rather than extracting individual ones in order to capture inter-demonstration relationships. They train a Determinantal Point Process (DPP)-based retriever to match scores of the LLM output, obtaining the most likely demonstration set through maximum a-posteriori during inference.
2.2 Jointly Optimizing Instructions and ICE
Most of the above work on ICE selection, however, considers a prompt without any instruction. In a separate line of work, methods for instruction optimization have been proposed (Zhou et al., 2023; Pryzant et al., 2023), but these do not consider ICE in the prompt. A recent RL-based method Zhang et al. (2022a) optimizes over both discrete instruction and example sets in the prompt. However, the search space for unconstrained joint optimization becomes exponentially large for large instruction and example spaces.
While these works show the importance of optimizing instructions and examples for various tasks, they do not perform a comprehensive comparison of the two. In this paper, our objective is to compare the two task-learning paradigms---instruction following and in-context learning---and evaluate their relative importance across a diverse range of tasks. We find that the choice of optimal task-learning paradigm depends on the task, and propose a method to determine which of these optimizations is expected to benefit LLM performance.
3 Problem Formulation
Optimizing ICE amounts to selecting the best $k$ examples from a given pool of candidate examples that should be included in the prompt. However, optimizing the instruction is not a straightforward optimization because the space of possible instructions can be extremely large. Hence, we first provide a instruction template covering five types of instructions that we consider in this paper, and then define our research questions.
3.1 Prompt Structure and Instruction Set
For a fair comparison across instructions and generalizability to new tasks, we use a standard prompt template consisting of an instruction, $k$-shot demonstrations $(x_{i},y_{i})$ from a training set $\mathcal{D_{T}}$, and a query $x$ from the test set with a cue for the model to predict $y$. We define a consistent structure to create a general set of instructions, iteratively adding useful, task-specific information to the instruction starting with no instruction and following a guideline similar to Zhang et al. (2023). A consistent template without human creativity ensures that our findings are not constrained to a narrow distribution of carefully-crafted prompts. The gradual addition of detailed information in the instruction serves as a proxy for "improved" instructions based on the results from Zhang et al. (2023).
  • No Instruction (NI): We provide no instruction, so LLM is expected to learn only from examples.
  • Task Definition with Label Space (+LS): Including the label space in the task definition. (e.g. [‘‘positive’’, ‘‘negative’’] for a binary sentiment classification task.
  • Delusive Instruction (DI): We also evaluate the model’s task learnability when the instruction is misleading, i.e., describing a task with shuffled labels. (e.g. ‘‘positive’’ for ‘‘negative’’).
Note that we presented the instructions in increasing order of detail or effectiveness, except the Delusive Instruction which is expected to have the lowest effectiveness. For generation tasks, since there is no notion of label space, we introduce an additional variant of detailed instruction:
  • Task Definition with Rules (+R): A simple task definition with syntactic rules for the task.
For generation tasks, we experiment with the ‘No-Instruction’ and ‘Task-Definition’ cases in the same way as defined for classification tasks. Some examples are given in Appendix A.1.
3.2 Research Questions
We focus on the following research questions.
RQ1: Do all tasks benefit from in-context example optimization?
To investigate whether ICE optimization is needed for a given task, we compare the LLM’s performance with different kinds of ICE. Specifically, we compare bins of ICE at different distances from a given query and check how the LLM’s performance varies. To understand how the effect of ICE optimization changes with the instruction quality, we repeat the experiment for the different instructions listed above. We evaluate various tasks based on the NICE metric defined in Eq. (2).
RQ2: How do ICE selection methods compare to random selection with a detailed instruction?
Here we directly compare ICE optimization methods versus random ICE selection and study how the LLM’s performance gap varies with different instruction types.
RQ3: Given a detailed instruction, does providing the correct labels for ICE matter? Past works have studied the impact of noisy Wei et al. (2023) and biased Gupta et al. (2023) labels on LLM performance. We study how those conclusions may not hold in the presence of detailed instructions.
Min et al. (2022) identify four aspects of demonstrations that serve as a learning signal to the LLM: the input-label mapping, the distribution of the input text, the label space and the format. Concretely, they say that in-context examples play a role by informing the model of the label space or the format of the output. Pan et al. (2023) discuss the phenomena of task learning and task recognition as the two ways through which an LLM learns from demonstrations by comparing the performance of randomly labeled examples against abstract and gold-labeled ones.
Based on these works, we evaluate the role of input-label mapping for in-context learning, under the presence of a detailed task instruction. Specifically, we investigate whether in-context examples are being used for information about the input-label mapping or they are simply serving as information for task recognition: providing the input distribution, label space or output format. To this extent, we introduce perturbations in the in-context examples by randomizing their ground-truth labels. In the case of classification tasks, we assign randomly chosen labels $y\sim\mathcal{Y_{T}}$ to the inputs of the in-context examples, where $\mathcal{Y_{T}}$ denotes the label space of the output for a particular task $\mathcal{T}$. In the case of generation tasks, we shuffle the ground truth values $y$ of the ICE amongst themselves to introduce a perturbation.
@@ 3.2 Research Questions
4 NICE: Measuring Invariability to ICE
In addition to answering the research questions above, we provide a practical metric to assess the effectiveness of optimizing ICE. That is, to decide whether to optimize instruction or in-context examples, we develop a task-specific metric for a given instruction that measures the invariability of task performance to ICE selection.
We assume that each task is associated with a training set from which ICE are selected for a given query. Given a new query, we assume access to a scoring function $f$ of the candidate examples and the query. To study invariance to ICE selection, for each query in the test set, we partition available training data of candidate examples using the function $f$ into multiple bins. Subsequently, for each bin, given a test sample of queries, we compute the model’s performance using a pre-specified instruction $I$ and random ICE from the bin for each query (and then averaging the resultant accuracy). This yields the score $S$ for a bin,
@@ 4 NICE: Measuring Invariability to ICE
where $Y_{LM}$ is the performance measure (higher is better) of the language model, $\mathcal{D}$ is the set of test queries, I is the given instruction and $b^{x}_{j}\in\mathcal{B}$ is the $j$-th bin for query $x$. Intuitively, given a task, high variance in the scores $S$ across bins indicates a task where ICE selection may be important.
@@ 4.1 Metric Requirements
Before proposing a metric, we first outline some properties to be followed by the proposed metric so that it can capture returns of ICE optimization for a given task and a given instruction.
@@
  • Property 3: If $S(j,\text{I},\mathcal{D})\approx\max\limits_{j}S(j,\text{I},\mathcal{D})$ for all $j$, i.e., the performance of the model is invariant to the bin from which the examples are chosen, the metric should be close to $1$.
  • Property 4: Let $b_{\text{max}}=\arg\max_{j}S(j,\text{I},\mathcal{D})$ be the best bin. If, for all $j\neq b_{\text{max}}$, $S(j,\text{I},\mathcal{D})\ll S(b_{\text{max}},\text{I},\mathcal{D})$, this represents a strictly retrieval task where the budget needs to be spent on ICE optimization. In this scenario, the metric should approach 00.
Since API calls to LLMs are both costly and time consuming, another requirement is that the metric should be computable much more efficiently than typical ways to optimize over the search space of all candidate examples. In the next section, we introduce the NICE metric and in Appendix A.2, we prove that it follows all Properties 1-4.